OpenStudy (anonymous):

Estimate 13.96 x 6.2

OpenStudy (anonymous):

A. about 84 B. about 70 C. about 98 D. about 78

OpenStudy (mathstudent55):

I know that 14 * 3 = 42 That means 14 * 6 = 82 Since 6.2 is a little more than 6, I estimate 85

OpenStudy (anonymous):

should i just put 84 since its the closest?

OpenStudy (phi):

they might be thinking 13.96 is about 14 6.2 is about 6 6*14= 6*10 + 6*4 = 60+24= 84

OpenStudy (anonymous):

now i understand. thankyou so muchhhh !!

OpenStudy (mathstudent55):

Yes, use 84.